OpenStudy (anonymous):

I have a separable differential equation where (dy/dx)=24x^3y^5, and the initial value is at (-1, 1). Is the right start (-1/5)ln(y^5)= -30x^4+C?

OpenStudy (anonymous):

\[dy/dx=24x^3y^5\] \[\rightarrow dy/y^5=24x^3dx\] \[\int\limits_{?}^{?}dy/y^5=\int\limits_{?}^{?}24x^3dx\] \[\rightarrow -y^4/4=6x^4 +c\] Then solve it given the initial conditions

OpenStudy (anonymous):

awesome nadeem, thanks for the quick reply :D

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

so once I have c equal to 24, I just solve the original equation for y?

OpenStudy (anonymous):

or solve from -y^4/4=6x^4+24?

OpenStudy (anonymous):

solve from -y^4/4=6x^4+24 by plugging in y=1 and x=-1

OpenStudy (anonymous):

I thought plugging in the initial values gave me a method to solve c? but once I solve c and have to solve the equation for y do I solve from -y^4/4=6x^4+24?

OpenStudy (anonymous):

yeah, sorry for the confusion

OpenStudy (anonymous):

no problem at all, thanks for all your help! you made my night of homework a lot less frustrating! :D

OpenStudy (anonymous):

I'm working on differential equations as well.... so frustrating sounds just about right